OpenStudy (anonymous):

What is the value of the expression? 4!(7 – 4)!

OpenStudy (anonymous):

Don't Give Me The Anwser ... Help Me Understand How To Work It Out.

OpenStudy (anonymous):

4! = ? (7-4)! = 3! = ?

OpenStudy (anonymous):

4! times 7 = 168 and 4! times4 = 69 and 168 -69 = ?

OpenStudy (anonymous):

haha im not sure if im rite thou

OpenStudy (anonymous):

Is Tht A Negative Or Minus Sign -69?

OpenStudy (anonymous):

It should be: 4!(3!) = 24(6) = 144

OpenStudy (anonymous):

Thank You ^.^